solve using the substitution method r-7s=0 and 9r-2s=61
Answer by Alan3354(69443)

solve using the substitution method r-7s=0 and 9r-2s=61
-------------
r-7s=0 --> r = 7s Sub for r in the other equation.
9r-2s=61